I'm having trouble solving this problem:
David and Daniel's father gave them equal amounts of money. David earned money doing chores until he had twice as much as he started with. Daniel didn't earn any more, but he received $12.00 for his birthday. If the boys still have the same amount of money, how much did their father give each one? I also need help setting up this problem. Thank you.

Answer by fcabanski(1391)

They both started out with x, the unknown amount of money.


David ended up with twice the original amount, which is 2x


Daniel had the original x plus $12 from his birthday. So he ended up with x+12.


Daniel and David ended up with equal amounts, so 2x = 12+x.


Subtract x from both sides. x=12.


Check:2x = 12+x


2*12 = 12 + 12 = 24.